OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution I got D is that correct ?

OpenStudy (anonymous):

@Hero @BangkokGarrett

OpenStudy (anonymous):

Anyone ?

OpenStudy (anonymous):

I got B - how did you get d?

OpenStudy (anonymous):

I just couldn't find the correct answer, could you show me how you got B ?

OpenStudy (anonymous):

\[−3|2x + 6| = −12\]Isolate the absolute value by dividing both sides by -3\[\left| 2x+6 \right|=4\] To deal with the absolute value, just we have the equation as is, and the opposite of it, just without the absolute value sign. \[2x+6=4 \] \[2x+6=-4\] Just solve for x for both of those to get the answer

OpenStudy (anonymous):

Okay can you explain one more for me ?

OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

I got A on that one^

OpenStudy (anonymous):

Check your work